OpenStudy (anonymous):

Need help, Find and simplify the difference quotient f(x+h)-f(x)/h, h≠0 for given function. f(x)=1/5x

OpenStudy (phi):

Did you try to write down the difference? if it is \( f(x) = \frac{1}{5x} \) you would do this \[ \frac{ \frac{1}{5(x-h)}- \frac{1}{5x} }{h} \\ = \frac{1}{h} \left( \frac{1}{5(x-h)}- \frac{1}{5x}\right) \] I would put the two fractions over a common denominator as the next step. Can you try?